# Ferroelectric switchable altermagnetism

### v1

@@ Abstract
Recently, altermagnetism [13, 14] has garnered considerable attention as a promising avenue for achieving novel spintronic properties. Altermagnets are a class of collinear antiferromagnetic (AFM) materials characterized by alternating spin polarization across reciprocal space due to breaking of certain spatial symmetries, even though they have no net macroscopic magnetization. The order parameter S featuring altermagnetism can be represented by the energy splitting between the two spin channels in certain paths of the Brillouin zone $\Delta E_{k}^{s}=E_{k}^{\uparrow}-E_{k}^{\downarrow}$. The spin-split bands of Bloch electrons in altermagnets create the possibility of designing new types of spintronic devices, e.g., spin-filtering magnetic tunnel junctions [15, 16]. Due to the combination of the advantages of FM and AFM in terms of spin splitting, altermagnetism is also considered a third type of collinear magnetism [13, 17].
In this Letter, we introduce the ferroelectric switchable altermagnets, a new class of materials which shows not only the coexistence of ferroelectricity and altermagnetism, but also the synergistic coupling of the two order parameters, leading to a reversal of spin-character in the splitted energy bands by switching the ferroelectric polarization. As illustrated in Fig. 1(a), in specific circumstances, the altermagnetic spin degree of freedom is tightly correlated to the electric polarization, indicating a distinct form of magnetoelectric coupling. Such a mechanism enables the control of a nonvolatile spin-split electronic dispersion through the external electric field, providing an efficient route to achieve robust all-in-one multiferroic memory devices. For example, Fig. 1(b) illustrates an electric-field-controllable spin-filtering altermagnetic tunnel junction device without the need of switching the Néel vector. By considering whether the spin-polarized Fermi surface between the sample and the substrate match or not [13], the resistance can be switched between high and low by simply switching the ferroelectric polarization. By applying the state-of-the-art spin-group symmetry analyses, we searched through all the MAGNDATA database [18] and screened 22 ferroelectric altermagnets, in which only two candidates support the synergistic switching of altermagnetic spin splitting through ferroelectric polarization.
@@
Figure 1: (a) Schematic illustration of ferroelectric switchable altermagnetism, where the altermagnetic spin splitting S strongly couples to the ferroelectric polarization P. (b) A non-volatile spin filtering tunnel junction device design with ferroelectric switchable altermagnets.
@@
In general, the coexistence of electric polarization P and altermagnetic spin splitting $\Delta E_{k}^{S}$ (denoted by S) does not guarantee their coupling. Therefore, we next focus on the search of symmetry operations that synergistically switch P and S, and if these operations correspond to reasonable flipping paths under an external electric field, which does not change the Néel order. For example, considering an initial state $\psi_{k}^{i}(P,S)$, one can always flip both electric and spin polarization by applying a joint operation $\mathcal{PT}$ to the system to get a reversed state $\mathcal{PT}\psi_{k}^{i}(P,S)=\psi_{k}^{f}(-P,-S)$. However, in the transition path connecting $\psi_{k}^{f}(-P,-S)$ and $\psi_{k}^{i}(P,S)$ phases, the Néel vector may also be reversed, indicating a transition path that cannot be achieved by electric field only. By going through the list in TABLE SI in Supplementary Materials, we found that the sign of $\Delta E_{k}^{S}$ for most candidates remains unchanged during the polarization reversal, as exemplified by PbNiO3 in Supplementary Materials. Interestingly, there are two hybrid improper ferroelectric materials, Ca3Mn2O7 and [C(NH2)3]Cr(HCOO)3, as prototypical candidates of ferroelectric switchable altermagnets.
@@
Altermagnetism in hybrid improper ferroelectric Cr-MOF - Hybrid improper ferroelectricity arises from a combination of two or more lattice modes that individually preserve inversion symmetry but collectively breaks it, resulting in a spontaneous polarization. Several candidate materials from Class 1 in TABLE SI belong to this category, including Ca3Mn2O7 (SG $Cmc2_{1}$) [27] and the hybrid organic-inorganic perovskite [C(NH2)3]M(HCOO)3 (SG $Pna2_{1}$) with M being transition-metal elements Cu or Cr. Below we use Cr-MOF (see Fig. 2(a)) to illustrate the design principles of ferroelectric switchable altermagnets for two reasons: (a) it exhibits the largest spin splitting near the Fermi level among this material family, and (b) various AFM configurations can be realized in Cr-MOF through compressive strain, offering an additional degree of freedom for phase manipulation [28]. Moreover, the compositional tunability in this family offers more flexibility in designing materials for spin-splitting magnitude optimization. While A-, C- and G-type AFM Cr-MOF are all altermagnets, here we focus on the C-type configuration, as it manifests the largest spin splitting near the Fermi level. Further discussions on other MOFs and AFM configurations can be found in the Supplementary Materials.
The spin-polarized band structure obtained from density functional theory (DFT) calculations is shown in Fig. 2(b). The band gap is about 3 eV, ensuring a good insulating background that is favorable for ferroelectrics. Notably, a large spin splitting is observed at the valence band maximum (VBM), with the largest energy difference being approximately 20 meV. Such spin splitting occurs at the relatively low-symmetry points in reciprocal space, such as along the $\Gamma(0,0,0)\rightarrow R(0.5,0.5,0.5)$ path. The spin-polarized isoenergy contour at $E=-0.05$ eV is shown in Fig. 2(c), showing a deviated feature compared with the so-called “d-wave” altermagnetism because of the absence of four-fold symmetry. Furthermore, the sign of $\Delta E_{k}^{S}$, a well as the spin polarization of a given spin-polarized band, are odd under mirror operations {1||m100} and {1||m010}, and even under {1||m001}, as illustrated in Fig. 2(e).
@@
Synergistically switch of electric and spin polarizations - Here, a state with the coexistence of polarization $P$ ($P//z$) and altermagnetism $S$ is denoted as the $(P,S)$ state. The SSG operations that map $P\rightarrow-P$ and $S\rightarrow-S$ are analyzed and listed in Fig. S3., While the SSG operations connecting $P$ state to $-P$ state can be one of the following: {1||-1}, {1||2100}, {1||2010}, {-1||-1}, {-1||2100}, and {-1||2010}, only three of them map the $(P,S)$state to the $(-P,-S)$ state, i.e., {1||2100}, {1||2010}, and {-1||-1}. In addition, whether the $(-P,-S)$ state can be experimentally realized upon the application of an electric field, however, depends on the energy barrier between the two phases, which is ultimately governed by the internal atomic displacements during the phase transition.
The search for the reversed polarization state can also be achieved by considering the mode decomposition of hybrid improper ferroelectricity. The electric polarization of Cr-MOF along the z direction is activated by two non-polar lattice distortion modes through a trilinear coupling $F=\gamma Q_{X_{1}^{-}}Q_{X_{4}^{+}}Q_{P}$, where $Q_{X_{1}^{-}}$, $Q_{X_{4}^{+}}$ and $Q_{P}$ represent the amplitudes of the two zone-boundary non-polar distortion modes and the polar distortion mode, respectively, relative to the high symmetry Imma phase [19]. Reversing either of the zone-boundary modes flips the polarization $P\rightarrow-P$, while, reversing both, preserves $P$. As shown in Fig. 3(a), the $X_{1}^{-}$ mode consist of two major distortions: a “scissor”-like distortion for the equatorial Cr-O bonds and a rotation of the guanidinium molecule by about 42°. The $X_{4}^{+}$ mode corresponds primarily to a Q2 Jahn-Teller (JT) distortion of the Cr octahedra, with minimal impact on the organic molecules. Starting from the Imma structure, imposing only the $X_{1}^{-}$ ($X_{4}^{+}$) mode results in a centrosymmetric Pnna (Pnma) structure. The combination of both $X_{1}^{-}$ and $X_{4}^{+}$ modes leads to the Pna21 structure, a polar SG that induces the softening of the polar distortion. Since the polar distortion is a secondary effect arising from the coupling of the zone-boundary modes, the interplay between these two modes fully determines the final structure and its related properties in this system. For simplicity, and without loss of generality, we will use these two unstable modes ($X_{1}^{-}$ and $X_{4}^{+}$) as variables in the discussion below.
@@
The spin reversal of the isoenergy contour for the $(-P,-S)$ state shown in Fig. 2(d) strongly supports the possibility to control altermagnetic spin splitting by flipping the ferroelectric polarization in this system. To quantify this transition, we define a normalized interpolating structural parameter $\xi$, where $\xi=1$ represents the $(P,S)$ structure and $\xi=-1$ corresponds to either the $(-P,S)$ or $(-P,-S)$ structure. The evolution of $\Delta E_{k}^{s}$ during the structural transition is plotted in Fig. 3(d). We found that $\Delta E_{k}^{s}$ changes monotonically during the $(P,S)\rightarrow(-P,-S)$ transition, while in the $(P,S)\rightarrow(-P,S)$ transition it reduces to zero at the Pnma phase and then regresses to its original value. Such behavior arises because $\Delta E_{k}^{s}$ is odd under {1||2100} but even under {1||m001}, consistent with the symmetry presented in Fig. 2(e). Furthermore, the microscopic mechanism of the reversal of the altermagnetic spin splitting by reversing the $X_{4}^{+}$ mode can be understood by the change of the Q2 JT distortion pattern, as illustrated in the insets of Fig. 3(d). Since the JT distortion is known to couple with the orbital ordering in oxide perovskites [29] and MOF systems [30], our results provide a concrete example of how altermagnetism can be entangled with orbital ordering [31], offering new insights into the control of spin splitting through orbital as well as lattice degrees of freedoms.
@@
Conclusion and discussion - Recently, the gate-field control of different spin channels in a layered altermagnet through spin-valley-layer locking effect has been proposed [36], highlighting the importance for the manipulation of spin degree of freedom in altermagnets. However, such a non-versatile strategy requires the constant application of an external electric field, which is not favorable for storage devices unless further integration with other materials. On the other hand, sliding ferroelectricity can also be used to tune the $\mathcal{PT}$ symmetry in a complex van der Waals sandwich heterojunction systems [37], which eventually controls emergence of the altermagnetic state. However, it requires an intrinsic AFM monolayer system in the center of the heterojunction, which is somehow rare. Considering that the hybrid improper ferroelectric materials can be artificially designed by chemistry or strain engineering [38], our study provides a wide configuration space for designing versatile ferroelectric switchable altermagnets.
Our approach provides a novel method for controlling spin properties through ferroelectric polarization, which differs from traditional magnetoelectric multiferroics where ferroelectricity is used to control the net magnetic moment. For class 1 candidates, such as Ca3Mn2O7, both approaches can work synergistically, offering multiple degrees of freedom for material manipulation. The coupling of electric polarization to both the weak FM moment and spin splitting provides enhanced versatility in device design. In contrast, for class 2 candidates where there is no net magnetic moment and most traditional magnetic detection method such as magneto-optical Kerr effect [39] would be ineffective, our approach still offers a viable pathway to achieve magnetoelectric coupling by correlating the electric polarization with spin splitting in reciprocal space. To conclude, ferroelectric switchable altermagnetism is a new type of magnetoelectric coupling beyond what is achievable in traditional multiferroics based solely on the spin-orbit interaction, paving an new functional materials platform for designing all-electric-controlled spintronic devices.
@@ Acknowledgements.
A.S. thanks useful discussions with P. Radaelli (Oxford University), P. Barone (CNR-SPIN), Sang-Wook Cheong (Rutgers University) and Xinfeng Chen (Xi’an Jiaotong University). This work was supported by the National Key R&D Program of China under Grant No. 2019YFA0704900, the National Natural Science Foundation of China under Grant No. 12274194, Guangdong Provincial Key Laboratory for Computational Science and Material Design under Grant No. 2019B030301001, Shenzhen Science and Technology Program (Grants No. RCJC20221008092722009 and No. 20231117091158001), the Innovative Team of General Higher Educational Institutes in Guangdong Province (Grant No. 2020KCXTD001), Guangdong Provincial Quantum Science Strategic Initiative under Grant No. GDZX2401002, the Science Technology and Innovation Commission of Shenzhen Municipality (JCYJ20210324104812034), Natural Science Foundation of Guangdong Province (2021A1515110389) and Center for Computational Science and Engineering of Southern University of Science and Technology. This work has been funded by the European Union - NextGenerationEU, Mission 4, Component 1, under the Italian Ministry of University and Research (MUR) National Innovation Ecosystem grant ECS00000041-VITALITY-CUP B43C22000470005. We also acknowledge the support by bilateral agreement for scientific collaboration CNR and NSFC (China), ”Ferroelectric and chiral hybrid organic inorganic perovskites”, for the years 2024-2025 (CUP B53C2300716000). K.Y. was supported by an individual grant (No. CG092501) at Korea Institute for Advanced Study.

### v2

@@ Abstract
Recently, altermagnetism has garnered considerable attention [13, 14, 15, 16] as a promising avenue for achieving novel spintronic properties. Altermagnets are a class of collinear antiferromagnetic (AFM) materials characterized by alternating spin polarization across reciprocal space due to breaking of certain spatial symmetries, even though they have no net macroscopic magnetization. The order parameter S featuring altermagnetism can be represented by the energy splitting between the two spin channels in certain paths of the Brillouin zone $\Delta E_{k}^{s}=E_{k}^{\uparrow}-E_{k}^{\downarrow}$. The spin-split bands of Bloch electrons in altermagnets create the possibility of designing new types of spintronic devices, e.g., spin-filtering magnetic tunnel junctions [17, 18]. Due to the combination of the advantages of FM and AFM in terms of spin splitting, altermagnetism is also considered a third type of collinear magnetism [14, 19].
In this Letter, we introduce the ferroelectric switchable altermagnets, a new class of materials which shows not only the coexistence of ferroelectricity and altermagnetism, but also the synergistic coupling of the two order parameters, leading to a reversal of spin-character in the splitted energy bands by switching the ferroelectric polarization. As illustrated in Fig. 1(a), in specific circumstances, the altermagnetic spin degree of freedom is tightly correlated to the electric polarization, indicating a distinct form of magnetoelectric coupling. Such a mechanism enables the control of a nonvolatile spin-split electronic dispersion through the external electric field, providing an efficient route to achieve robust all-in-one multiferroic memory devices. For example, Fig. 1(b) illustrates an electric-field-controllable spin-filtering altermagnetic tunnel junction device without the need of switching the Néel vector. By considering whether the spin-polarized Fermi surface between the free layer and the fixed layer match or not [14], the conductance $\sigma$ can be switched between high and low by simply switching the ferroelectric polarization. By applying the state-of-the-art spin-group symmetry analyses, we searched through all the MAGNDATA database [20] and screened out 22 ferroelectric altermagnets, in which only two candidates support the synergistic switching of altermagnetic spin splitting through ferroelectric polarization.
@@
Figure 1: (a) Schematic illustration of ferroelectric switchable altermagnetism, where the altermagnetic spin splitting S strongly couples to the ferroelectric polarization P. (b) A non-volatile spin filtering tunnel junction device design with ferroelectric switchable altermagnets. The Fermi surfaces of the fixed layer and the free layer are also illustrated. The electric polarization $P$ can be used to control the spin polarization of the free layer, leading to a transition between high- and low-conductance ($\sigma$) states.
@@
In general, the coexistence of electric polarization P and altermagnetic spin splitting $\Delta E_{k}^{S}$ (denoted by S) does not guarantee their coupling. Therefore, we next focus on the search of symmetry operations that synergistically switch P and S, and if these operations correspond to reasonable flipping paths under an external electric field, which does not change the Néel order. For example, considering an initial state $\psi_{k}^{i}(P,S)$, one can always flip both electric and spin polarization by applying a joint operation $\mathcal{PT}$ to the system to get a reversed state $\mathcal{PT}\psi_{k}^{i}(P,S)=\psi_{k}^{f}(-P,-S)$. However, in the transition path connecting $\psi_{k}^{f}(-P,-S)$ and $\psi_{k}^{i}(P,S)$ phases, the Néel vector may also be reversed, indicating a transition path that cannot be achieved by electric field only. By going through the list in TABLE SI in Supplementary Materials, we found that the sign of $\Delta E_{k}^{S}$ for most candidates remains unchanged during the polarization reversal, as exemplified by PbNiO3 in Supplementary Materials Sec. III. Interestingly, there are two hybrid improper ferroelectric materials, Ca3Mn2O7 and [C(NH2)3]Cr(HCOO)3, as prototypical candidates of ferroelectric switchable altermagnets.
@@
Altermagnetism in hybrid improper ferroelectric Cr-MOF - Hybrid improper ferroelectricity arises from a combination of two or more lattice modes that individually preserve inversion symmetry but collectively breaks it, resulting in a spontaneous polarization. Several candidate materials from Class 1 in TABLE SI belong to this category, including Ca3Mn2O7 (SG $Cmc2_{1}$) [29] and the hybrid organic-inorganic perovskite [C(NH2)3]M(HCOO)3 (SG $Pna2_{1}$) with M being transition-metal elements Cu or Cr. Below we use Cr-MOF (see Fig. 2(a)) to illustrate the design principles of ferroelectric switchable altermagnets for two reasons: (a) it exhibits the largest spin splitting near the Fermi level among this material family, and (b) various AFM configurations can be realized in Cr-MOF through compressive strain, offering an additional degree of freedom for phase manipulation [30]. Moreover, the compositional tunability in this family offers more flexibility in designing materials for spin-splitting magnitude optimization. While A-, C- and G-type AFM Cr-MOF are all altermagnets, here we focus on the C-type configuration, as it manifests the largest spin splitting near the Fermi level. Further discussions on other MOFs and AFM configurations can be found in the Supplementary Materials Sec. IV.
The spin-polarized band structure obtained from density functional theory (DFT, see Supplementary Materials Sec. I for details) calculations is shown in Fig. 2(b). The band gap is about 3 eV, ensuring a good insulating background that is favorable for ferroelectrics. Notably, a large spin splitting is observed at the valence band maximum (VBM), with the largest energy difference being approximately 20 meV. Such spin splitting occurs at the relatively low-symmetry points in reciprocal space, such as along the $\Gamma(0,0,0)\rightarrow R(0.5,0.5,0.5)$ path. The spin-polarized isoenergy contour at $E=-0.05$ eV is shown in Fig. 2(c), showing a deviated feature compared with the so-called “d-wave” altermagnetism because of the absence of four-fold symmetry. Furthermore, the sign of $\Delta E_{k}^{S}$, a well as the spin polarization of a given spin-polarized band, are odd under mirror operations {1||m100} and {1||m010}, and even under {1||m001}, as illustrated in Fig. 2(e).
@@
Synergistically switch of electric and spin polarizations - Here, a state with the coexistence of polarization $P$ ($P//z$) and altermagnetism $S$ is denoted as the $(P,S)$ state. The SSG operations that map $P\rightarrow-P$ and $S\rightarrow-S$ are analyzed and listed in Fig. S3. While the SSG operations connecting $P$ to $-P$ state can be one of the following: {1||-1}, {1||2100}, {1||2010}, {-1||-1}, {-1||2100}, and {-1||2010}, only three of them map the $(P,S)$state to the $(-P,-S)$ state, i.e., {1||2100}, {1||2010}, and {-1||-1}. In addition, whether the $(-P,-S)$ state can be experimentally realized upon the application of an electric field, however, depends on the energy barrier between the two phases, which is ultimately governed by the internal atomic displacements during the phase transition.
The search for the reversed polarization state can also be achieved by considering the mode decomposition of hybrid improper ferroelectricity. The electric polarization of Cr-MOF along the z direction is activated by two non-polar lattice distortion modes through a trilinear coupling $F=\gamma Q_{X_{1}^{-}}Q_{X_{4}^{+}}Q_{P}$, where $Q_{X_{1}^{-}}$, $Q_{X_{4}^{+}}$ and $Q_{P}$ represent the amplitudes of the two zone-boundary non-polar distortion modes and the polar distortion mode, respectively, relative to the high symmetry Imma phase [21]. Reversing either of the zone-boundary modes flips the polarization $P\rightarrow-P$, while, reversing both preserves $P$. As shown in Fig. 3(a), the $X_{1}^{-}$ mode consists of two major distortions: a “scissor”-like distortion for the equatorial Cr-O bonds and a rotation of the guanidinium molecule by about 42°. The $X_{4}^{+}$ mode corresponds primarily to a Q2 Jahn-Teller (JT) distortion of the Cr octahedra, with minimal impact on the organic molecules. Starting from the Imma structure, imposing only the $X_{1}^{-}$ ($X_{4}^{+}$) mode results in a centrosymmetric Pnna (Pnma) structure. The combination of both $X_{1}^{-}$ and $X_{4}^{+}$ modes leads to the Pna21 structure, a polar SG that induces the softening of the polar distortion. Since the polar distortion is a secondary effect arising from the coupling of the zone-boundary modes, the interplay between these two modes fully determines the final structure and its related properties in this system. For simplicity, and without loss of generality, we will use these two unstable modes ($X_{1}^{-}$ and $X_{4}^{+}$) as variables in the discussion below.
@@
The spin reversal of the isoenergy contour for the $(-P,-S)$ state shown in Fig. 2(d) strongly supports the possibility to control altermagnetic spin splitting by flipping the ferroelectric polarization in this system. To quantify this transition, we define a normalized interpolating structural parameter $\xi$, where $\xi=1$ represents the $(P,S)$ structure and $\xi=-1$ corresponds to either the $(-P,S)$ or $(-P,-S)$ structure. The evolution of $\Delta E_{k}^{s}$ during the structural transition is plotted in Fig. 3(d). We found that $\Delta E_{k}^{s}$ changes monotonically during the $(P,S)\rightarrow(-P,-S)$ transition, while in the $(P,S)\rightarrow(-P,S)$ transition it reduces to zero at the Pnma phase and then regresses to its original value. Such behavior arises because $\Delta E_{k}^{s}$ is odd under {1||2100} but even under {1||m001}, consistent with the symmetry presented in Fig. 2(e). Furthermore, the microscopic mechanism of the reversal of the altermagnetic spin splitting by reversing the $X_{4}^{+}$ mode can be understood by the change of the Q2 JT distortion pattern, as illustrated in the insets of Fig. 3(d). Since the JT distortion is known to couple with the orbital ordering in oxide perovskites [31] and MOF systems [32], our results provide a concrete example of how altermagnetism can be entangled with orbital ordering [33], offering new insights into the control of spin splitting through orbital as well as lattice degrees of freedom.
To illustrate the electric-field-controllable altermagnetic tunnel junction device, we consider a prototypical homojunction system proposed in Fig. 1(b), where both the fixed and free layers are composed of Cr-MOF. We evaluate the giant magnetoresistance (GMR) by the ratio between the DFT-calculated longitudinal conductivity of the spin-up and spin-down channels along the [110] direction [34]. Based on a 184-atom $\sqrt{2}\times\sqrt{2}$ supercell, our results suggest that the maximum GMR rate for this system could reach approximately 53% (see Fig. S5 and Supplementary Materials Sec. VI for details), validating the potential of ferroelectric switchable altermagnetism for non-volatile spintronic devices.
@@
Conclusion and discussion - Recently, the gate-field control of different spin channels in layered altermagnets through spin-valley-layer locking effect has been proposed [39, 40], highlighting the importance for the manipulation of spin degree of freedom in altermagnets. However, such a non-versatile strategy requires the constant application of an external electric field, which is not favorable for storage devices unless further integration with other materials. Sliding ferroelectricity [41] or phase transition between ferroelectric and antiferroelectric states [42] can also be used to tune the $\mathcal{PT}$ symmetry in the two dimensional van der Waals systems, which eventually control the emergence of the altermagnetic state. Considering that the hybrid improper ferroelectric materials can be artificially designed by chemistry or strain engineering [43], our study provides a wide configuration space for designing versatile ferroelectric switchable altermagnets.
Our approach provides a novel method for controlling spin properties through ferroelectric polarization, which differs from traditional magnetoelectric multiferroics where ferroelectricity is used to control the net magnetic moment. For the Class 1 candidate Ca3Mn2O7, the energy splitting ($\sim$ 100 meV) is even larger than that of Cr-MOF. Spin canting is allowed in this system, leading to a weak net ferromagnetic moment $M$. Our symmetry analysis shows that the order parameter for spin splitting $S$ follows the same transformation rules as that for $M$ [11], which can be detected by the anomalous Hall effect. Therefore, an additional order parameter $M$ is added to the coupled degrees of freedom, i.e., $(P,M,S)$, offering multiple tuning knobs for material manipulation and providing enhanced versatility in device design (see Supplementary Materials Sec. V for details). In contrast, for the Class 2 candidates with no net magnetic moment, the most traditional magnetic detection method such as magneto-optical Kerr effect [44, 16] could be effective, our approach still offers a viable pathway to achieve magnetoelectric coupling by correlating the electric polarization with spin splitting in reciprocal space. To conclude, ferroelectric switchable altermagnetism is a new type of magnetoelectric coupling beyond what is achievable in traditional multiferroics based solely on the spin-orbit interaction, paving a new functional materials platform for designing all-electric-controlled spintronic devices.
Note added - Recently, we became aware of a related study on Ca3Mn2O7 for the “altermagnetoelectric effect” [45], a concept similar to the ferroelectric switchable altermagnetism proposed in our work.
@@ Acknowledgements.
A.S. thanks useful discussions with P. Radaelli (Oxford University), P. Barone (CNR-SPIN), Sang-Wook Cheong (Rutgers University) and Xinfeng Chen (Xi’an Jiaotong University). This work was supported by the National Key R&D Program of China under Grant No. 2019YFA0704900, the National Natural Science Foundation of China under Grant No. 12274194 and 12474232, Guangdong Provincial Key Laboratory for Computational Science and Material Design under Grant No. 2019B030301001, Shenzhen Science and Technology Program (Grants No. RCJC20221008092722009 and No. 20231117091158001), the Innovative Team of General Higher Educational Institutes in Guangdong Province (Grant No. 2020KCXTD001), Guangdong Provincial Quantum Science Strategic Initiative under Grant No. GDZX2401002, the Science Technology and Innovation Commission of Shenzhen Municipality (JCYJ20210324104812034), Natural Science Foundation of Guangdong Province (2021A1515110389) and Center for Computational Science and Engineering of Southern University of Science and Technology. This work has been funded by the European Union - NextGenerationEU, Mission 4, Component 1, under the Italian Ministry of University and Research (MUR) National Innovation Ecosystem grant ECS00000041-VITALITY-CUP B43C22000470005. We also acknowledge the support by bilateral agreement for scientific collaboration CNR and NSFC (China), “Ferroelectric and chiral hybrid organic inorganic perovskites”, for the years 2024-2025 (CUP B53C2300716000). K.Y. was supported by an individual grant (No. CG092501) at Korea Institute for Advanced Study.